OpenStudy (anonymous):

how long would it take to stop a 0.145kg baseball traveling 45m/s using a force of 550N?

OpenStudy (anonymous):

f = m*a 550 = 0.145*a a = ?

OpenStudy (anonymous):

Kinematic equataion: v = v + at 0 = 45 + a*t solve for t, a is found from f=ma eqaution make sense?

OpenStudy (anonymous):

Answer=.01186 YES OR NO

OpenStudy (theeric):

Looks good to me.

OpenStudy (anonymous):

but if u apply back into the equation to check. It says it's not right

OpenStudy (theeric):

Hmm... \(F_\text{net}=ma\implies a=\dfrac{F_\text{net}}{m}\) Then, definition of acceleration, \(a=\dfrac{v_f-v_i}{\Delta t}\implies\dfrac{v_f-v_i}{a}=\Delta t\) And our variables are: \(t=?\) \(a=?\) \(v_f=0\) \(v_i=45\ [\text m/\text s^2]\) \(F = 550\ [\text N]\) \(m=.245\ [\text{kg}]\)

OpenStudy (anonymous):

mass is not .245

OpenStudy (theeric):

\(m=.145\ [\text{kg}]\), sorry! That's what I typed in. So, what do you get for \(a\)?

OpenStudy (theeric):

How are you checking it with an equation?

OpenStudy (anonymous):

@adamarbour "Answer=.01186 " "YES OR NO" yes, time to stop baseball = 0.01186 seconds. when you 'plugged it back in' did you plug it into the right equation? in the end, this is how the 2 equations should look. 550 = 0.145*3793 [True] 0 = 45 +(-3793 )*0.01186 [True]